# Comparing zero-shot self-explanations with human rationales in multilingual text classification

### v1

@@ Abstract
Instruction-tuned LLMs are able to provide an explanation about their output to users by generating self-explanations that do not require gradient computations or the application of possibly complex XAI methods. In this paper, we analyse whether this ability results in a good explanation by evaluating self-explanations in the form of input rationales with respect to their plausibility to humans as well as their faithfulness to models. For this, we apply two text classification tasks: sentiment classification and forced labour detection. Next to English, we further include Danish and Italian translations of the sentiment classification task and compare self-explanations to human annotations for all samples. To allow for direct comparisons, we also compute post-hoc feature attribution, i.e., layer-wise relevance propagation (LRP) and apply this pipeline to 4 LLMs (Llama2, Llama3, Mistral and Mixtral). Our results show that self-explanations align more closely with human annotations compared to LRP, while maintaining a comparable level of faithfulness.
@@
in multilingual text classification
Stephanie Brandlthanks:    Equal contribution. Department of Computer Science University of Copenhagen brandl@di.ku.dk                        Oliver Eberle Machine Learning Group Technische Universität Berlin oliver.eberle@tu-berlin.de
@@ 1 Introduction
Providing model explanations in order to increase trust and transparency into their decision making has been a key motivation for the field of Explainable AI (XAI) even before LLMs have found their way into everyday life. Nowadays, LLMs are being used for tasks like creative writing, help with homework but also for advice giving and translation while providing self-generated explanations at the same time.111www.washingtonpost.com/technology/2024/08/04/chatgpt-use-real-ai-chatbot-conversations This makes it all the more crucial to understand the quality of those self-explanations, how reliable they are and whether their faithfulness to the model and their plausibility to humans compare with other, widely investigated, post-hoc XAI methods.
@@
In this paper, we evaluate self-explanations from two text classification tasks for which human rationale annotations are available: sentiment classification and forced labour detection. We instruct 4 different LLMs (Mistral, Mixtral, Llama2 and Llama3) to solve the respective task and generate explanations based on the input text in a zero-shot experiment. We compare these with rationales provided on the same samples from various annotation studies and with state-of-the-art post-hoc explanations for Transformers, calculated based on the layer-wise relevance propagation (LRP) framework Ali et al. (2022) for each model respectively. For sentiment classification, we consider two different subsets from two different annotation studies, one also including Italian and Danish translations next to the English original. We evaluate on plausibility to the human annotations and faithfulness to the model decision, two established evaluation methods in the XAI literature DeYoung et al. (2020); Jacovi and Goldberg (2020). We further carry out qualitative analyses, for instance on the distribution of POS tags among the most frequently selected tokens, the differences between languages and the ability to follow precise instructions. We see this controlled study as a first step into a deeper understanding of the reliability and quality of self-explanations and we believe that by covering four language models, three languages and two very different domains of text classification tasks, with different levels of difficulty and text length we can provide valuable insights. We release our code on GitHub to enable reproducibility and potential future research.
@@ 2 Related Work
Generated self-explanations come with both new opportunities and challenges. Prior work in self-explanations for text has focused on new evaluation strategies and model improvements. Ye and Durrett (2022) evaluate whether including self-explanations can improve model performance on in-context learning while Madsen et al. (2024) proposed several instruction-based self-consistency checks to measure faithfulness in generated explanations.
Another line of work by Wiegreffe et al. (2022) seeks to improve free text self-explanations with the help of human-written explanations that are included in the instruction. Similarly to Kunz and Kuhlmann (2024), self-explanations are evaluated on a variety of properties by the means of human annotation. They are found to be generally true, grammatical and factual Wiegreffe et al. (2022) and further selective, to contain illustrative examples and rarely subjective according to Kunz and Kuhlmann. For those two papers GPT-3 on CommonsenseQA/NLI and GPT-4 on the Alpaca dataset were analysed, respectively. In our study, we consider human rationales as the ground truth for explaining a decision, against which we compare model self-explanations and post-hoc attributions.
@@
Our work focuses on a direct comparison of plausibility and faithfulness, using binary rationales and comparing them to post-hoc LRP attributions, which have been shown to faithfully reflect LLM predictions Ali et al. (2022); Achtibat et al. (2024). Additionally, we extend our analysis to a more complex text classification task–forced labor detection–and to multilingual settings, further broadening the scope and applicability of our study.
@@ 3.1 Datasets
SST We use two different subsets from the Stanford Sentiment Treebank (SST2, Socher et al. 2013) for binary sentiment classification on movie reviews. The first subset contains 263 samples from the validation and test split from SST2 with an average sentence length of 18 tokens. Human rationale annotations have been published for that subset by Thorn Jakobsen et al. (2023) where each sample has been annotated by multiple annotators, 8 on average, who were recruited via Prolific. Annotators were first asked to classify the sample into one of three classes: positive, neutral or negative where none of the sentences was assigned neutral as a gold label. In a second step, annotators should choose the parts of the input that support their label choice. We select the rationale annotations with the correct labels from the first step for further analysis. We averaged the binary rationales across all annotators (with correct label classification) and set a threshold of $0.5$ (after averaging) for the token selection. We additionally analyse the rationale annotations collected by Jørgensen et al. (2022) on a subset of 250 samples from the validation set of SST2. All samples were translated into Danish and Italian with an average sentence length of 15-17. Rationale annotation was carried out by 2 annotators per language (including English), who were native speakers with linguistic training. In contrast to the annotations collected by Thorn Jakobsen et al., the correct sentiment (positive or negative) was provided and the annotators were asked to select parts of the input that supported the gold label.
RaFoLa The authors of Mendez Guzman et al. (2022) published a Rationale-annotated corpus for Forced-Labour detection. This multi-class and multi-label dataset contains 989 English news articles that have been labelled and annotated according to 11 risk indicators defined by the International Labour Organization. Rationale annotations were carried out by two annotators who selected parts of the input to justify their label decision if they found evidence for any of the 11 labels. A subset of 100 articles was annotated by both annotators with a label agreement of $0.81$ (micro F1) and a rationale agreement of $0.73$ (intersection-over-union). The remaining articles were only annotated by one of the annotators. Each news article was assigned $1.2$ labels on average while $43\%$ were assigned with at least one label. For our analysis, we selected the 4 most frequent classes with occurrences between 117-256 out of the 989 articles. As we carry out zero-shot experiments on models that have not been fine-tuned on this task, we further convert this task into a binary classification task where we ask for a specific label once at a time. We are providing the definition of the respective forced labour indicator as part of the instruction, see Figure 9 & 10.
@@ 3.2 Rationale Extraction
For our experiments, we apply the same pipeline to the following 4 instruction fine-tuned LLMs: Llama2-13B,222meta-llama/Llama-2-13b-chat-hf Llama3.1-8B333meta-llama/Meta-Llama-3.1-8B-Instruct, Mistral-7B444mistralai/Mistral-7B-Instruct-v0.3 and quantized Mixtral-8x7B555mistralai/Mixtral-8x7B-Instruct-v0.1.
In a first step, we ask the model to classify the given text into positive/negative for SST and into yes/no depending on evidence for a specific risk indicator for the RaFoLa dataset. If the model manages to generate the correct answer, we ask it to generate rationales based on the relevant provided context of the input. In case of RaFoLa, we follow the original data collection and only request rationales if the respective risk indicator is present.
For the subsets in Italian and Danish, we have manually translated the prompts to the respective language with the help of native speakers.
3.3 LRP Relevances
To extract input attribution scores, we use layer-wise relevance propagation (LRP); a widely used and state-of-the-art XAI method to compute feature attributions Ali et al. (2022); Achtibat et al. (2024). Following the proposed propagation rules for Transformer models, we compute relevance scores for Llama and Mistral by backpropagating the logit for the correctly generated class token. While self-explanations and human annotations provide binary rationales, LRP assigns continuous scores to each token. To allow direct comparison, we used the number of tokens annotated by humans as the cut-off for LRP relevance scores in both tasks. Tokens were ranked by their scores, and the top-k tokens–where $k$ is the number of rationales from the human annotation–were selected for further analysis.
@@ 3.4 Constraining Self-Explanations
Initial experiments showed that without precise instructions, the model would return $80\%$ of the input tokens as rationales for SST where humans had annotated approximately $30\%$. This made comparisons difficult, so we chose to request a maximum number of tokens based on the number of annotated tokens by humans for each sample. Language models did not always follow this request but we could reduce the ratio of tokens to a comparable level with human annotations. For RaFoLa, this problem did not occur on the same level since the input texts were much longer and humans annotated entire phrases. We thus decided not to include an upper bound for the RaFoLa rationales. We will discuss ratios and instruction following in more details.
4 Results
We first show and discuss the main results of pair-wise agreement between the different kinds of rationales, i.e., plausibility scores, before further analysing their faithfulness, the differences between languages and a qualitative analysis.
@@ 4 Results
Figure 2: Pair-wise comparison scores (Cohen’s Kappa) between rationales on SST and multilingual SST (English, Danish and Italian). We compare human-annotated, model-generated, and post-hoc LRP rationales.
@@
#8 Physical and sexual violence
asdf
Figure 3: Pair-wise scores (Cohen’s Kappa) between rationales on RaFoLa for 4 different forced labour indicators.
@@ 4.1 Plausibility
We show pair-wise comparisons between human-annotated, model-generated and post-hoc computed rationales by calculating sample-wise Cohen’s Kappa scores between the binary scores and averaging across samples for different models. When considering human annotations as the ground truth, this is usually referred to as plausibility DeYoung et al. (2020). We here also show the comparison between post-hoc and self-generated rationales, i.e., self-explanations.
Cohen’s Kappa Cohen (1960) is a well-established method to measure inter-annotator agreement (IAA) between two annotators, in our case those are either the averaged human annotations or the two different types of model rationales (generated and post-hoc). We chose Kappa over F1 scores, which is also often used to evaluate IAA but comes with two obstacles. It is (i) driven by the imbalance of classes (here selected and not selected tokens) leading to a higher offset for annotations with a ratio of selected tokens closer to 0.5 and it (ii) does not consider randomness as a confounding factor. Both of these obstacles are taken into account when computing Cohen’s Kappa scores, leading to overall lower but more robust scores in comparison to F1. We report F1 scores in the Appendix in Section B.
@@
4.2 Task accuracy and rationale ratios
Table 1 presents model accuracies and ratios of identified rationale tokens for SST, multilingual SST (mSST), and RaFoLa datasets in humans and models. Accuracies for SST and mSST are generally high across models, while RaFoLa shows more variation, particularly lower accuracies and rationale ratios. The ratio of identified rationale tokens by humans tends to be higher for the shorter movie reviews in SST and mSST, and lower for the longer paragraphs in RaFoLa, suggesting that the human annotators have more sparsely identified relevant phrases when presented with longer texts. This can be attributed to the differences in samples across the datasets. While movie reviews in SST tend to be short and were written with the purpose of expressing a sentiment, the articles in RaFoLa were more descriptive of a specific situation or incident and might or might not include the violation of one (or more) of the 11 risk indicators. We further can assume that most models nowadays have seen the original English version of SST during training and are thus more familiar with this type of data.
Table 1: Model accuracies for SST, multilingual SST and RaFoLa, alongside the ratio of identified rationale tokens by humans and those of self-generated rationales across models.
acc ratio ratio model
llama llama3 mistral mixtral human llama llama3 mistral mixtral
SST 0.88 0.98 0.98 0.98 0.29 0.22 0.25 0.32 0.14
mSST (EN) 0.85 0.98 0.98 1.00 0.33 0.23 0.27 0.32 0.15
mSST (DK) 0.84 0.87 0.97 0.97 0.38 0.18 0.29 0.40 0.21
mSST (IT) 0.88 0.95 0.97 0.99 0.37 0.26 0.31 0.39 0.25
RaFoLA #1 0.50 0.47 0.65 0.39 0.05 0.05 0.23 0.12 0.16
RaFoLA #2 0.58 0.58 0.58 0.48 0.06 0.04 0.21 0.11 0.12
RaFoLA #5 0.88 0.92 0.89 0.81 0.04 0.02 0.08 0.06 0.05
RaFoLA #8 0.90 0.90 0.90 0.90 0.07 0.04 0.12 0.08 0.10
@@ 4.3 Faithfulness
Besides plausibility, faithfulness is the most commonly used evaluation approach to judge the quality of model explanations. Especially for feature attribution approaches, removal of most relevant features has been used to assess how faithful a feature subset is with respect to the model prediction, i.e. if removing a highly relevant subset will lead to a strong decrease of the prediction. We evaluated faithfulness by measuring the change in probability after masking the tokens as identified by the different rationales (human, self-generated and post-hoc).
@@
Compared to human and self-generated rationales, post-hoc attributions provide a relevance score for each token in the input prompt, requiring to binarize post-hoc attributions to allow for direct comparison as described in Section 3.3. Additionally, we included a baseline that randomly removes as many token as identified by humans. Our results for SST and RaFoLa are summarized in Figure 4. Average initial probability for the correct answer token is given as a dashed line for each model.
We find that levels of faithfulness for humans, generated, and post-hoc explanations are comparable, with human explanations being overall as faithful as those generated by models or post-hoc methods. For some cases, e.g. class #5 for RaFoLA, human rationales provide more faithful feature attributions than mistral explanations.
@@
Self-generated model explanations can be more faithful than post-hoc ones, and the reverse case can also be true, depending on the task and model. Overall, they provide similarly faithful model rationales. We further investigated the limited impact of removing features on the class token probability in the post-hoc setting and found that the most relevant tokens are typically part of the provided task instruction, such as the class definition or question. Contrastive explanations have been proposed to provide more task-specific attributions Yin and Neubig (2022) and we provide an additional analysis of the plausbility of post-hoc explanations in Section 4.8.
SST
mSST      English
Danish
Italian
RaFoLa        #1
#2
#5
#8
Figure 4: Faithfulness evaluation for SST and mSST (top row) and RaFoLa (bottom row). The probability after masking the tokens extracted from human rationale annotations, self-generated model rationales and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
4.4 Languages
From the set of models we consider for this study, all have been pre-trained on English data while Llama3 and Mixtral also have been pre-trained on Italian, but none of the model officially supports Danish. Considering this difference in language exposure, our results show that all models are able to solve the sentiment classification task (see Table 1) and are also able to extract meaningful rationales with plausibility on the same (or even superior) level than English. Plausibility scores between self-explanations and humans are ranging $0.46-0.74$ for Danish and Italian versus $0.4-0.69$ for English. Llama2 shows the biggest gap between English ($0.67$) and Danish/Italian ($0.46/0.52$) while Mixtral shows an increase in plausibility from English ($0.4$) to ($0.57$) for Danish. Llama3 shows high and stable plausibility scores across all languages ($\sim 0.7$). It seems all models can handle both Danish and Italian, whether this is based on data contamination during pre-training, e.g., training data for Llama2 has been reported to contain $0.11\%$ Italian Touvron et al. (2023), is unclear. Previous work has focused on zero-shot abilities in unseen languages but has mostly followed approaches with either multilingual models or translating test sets into English Ebrahimi and von der Wense (2024); Artetxe et al. (2023). To the best of our knowledge, there has not been any published work in a truly zero-shot fashion, we will leave that for future work.
Table 2: Most frequent tokens in the RaFoLa corpus(first row) and in rationales identified by human annotators, as well as self-generated and post-hoc explanations of Llama3 and Mistral for #1 and #8.
#1 Abuse of vulnerability #8 Physical and sexual violence

### v2

# Comparing zero-shot self-explanations with human rationales in text classification

@@ Abstract
Instruction-tuned LLMs are able to provide an explanation about their output to users by generating self-explanations. These do not require gradient computations or the application of possibly complex XAI methods. In this paper, we analyse whether this ability results in a good explanation. We evaluate self-explanations in the form of input rationales with respect to their plausibility to humans as well as their faithfulness to models. We study two text classification tasks: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. In addition to English, we include Danish and Italian translations of the sentiment classification task and compare self-explanations to human annotations for all samples. To allow for direct comparisons, we also compute post-hoc feature attribution, i.e., layer-wise relevance propagation (LRP) and analyse 4 LLMs. We show that self-explanations align more closely with human annotations compared to LRP, while maintaining a comparable level of faithfulness. This finding suggests that self-explanations indeed provide good explanations for text classification.
@@
in text classification
Stephanie Brandlthanks:    Equal contribution. Center for Social Data Science University of Copenhagen stephanie.brandl@sodas.ku.dk                        Oliver Eberle Machine Learning Group Technische Universität Berlin oliver.eberle@tu-berlin.de
@@ 1 Introduction
Providing model explanations to increase trust and transparency is a key motivation for the field of Explainable AI (XAI), with LLMs offering new ways to trace model decision-making. Nowadays, LLMs are being used for a wide range of tasks, ranging from creative writing and homework assistance to offering advice and translation, while providing self-generated explanations, i.e., self-explanations, in the process.111www.washingtonpost.com/technology/2024/08/04/chatgpt-use-real-ai-chatbot-conversations This makes it all the more important to understand the quality of those self-explanations, how reliable they are and whether their faithfulness to the model and their plausibility to humans compare with other, widely investigated, post-hoc XAI methods.
@@
In this paper, we evaluate self-explanations from two text classification tasks for which human rationale annotations are available: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. We instruct 4 different LLMs (Mistral, Mixtral, Llama2 and Llama3) to solve the respective task and generate explanations based on the input text in a zero-shot experiment. We compare these with rationales provided on the same samples from various annotation studies and with state-of-the-art post-hoc explanations for Transformers, calculated based on the layer-wise relevance propagation (LRP) framework Ali et al. (2022) for each model respectively, see Figure 1 for an example. For sentiment classification, we consider two different subsets from two different annotation studies, one also including Italian and Danish translations alongside the English original. Following established evaluation methods in the XAI literature DeYoung et al. (2020); Jacovi and Goldberg (2020), we assess the plausibility of model rationales by measuring their agreement with human annotations and evaluate faithfulness by determining the importance of selected tokens for the model’s decision. We extend our analysis by exploring differences across rationales and task settings, comparing distributions of POS tags, named entities, and frequently selected tokens. Our study takes an initial step toward a better understanding of the reliability and quality of self-explanations, for which we analyzed four language models, three languages, and two distinct text classification domains, varying in difficulty and text length. Our findings thus provide relevant insights for model interpretability and user trust in self-explanations, and we further support reproducibility and future research by openly releasing our code.
Contributions The main contributions of this work is (i) a controlled study to compare human annotations with model explanations generated by LLMs and computed post-hoc. (ii) We evaluate plausibility to humans, i.e., the level of agreement between model and human rationales and (iii) faithfulness, i.e., the relevance of selected rationale tokens for the task (model decision). (iv) We study two different text classification tasks: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. We include (v) Danish and Italian translations for the sentiment classification task alongside their English original. We further (vi) provide a qualitative analysis into the differences across languages/risk indicators with respect to frequent tokens and POS.
@@ 2 Related Work
Generated self-explanations present both new opportunities and challenges. Prior work in self-explanations for text has focused on new evaluation strategies and model improvements. Ye and Durrett (2022) evaluate whether including self-explanations can improve model performance on in-context learning while Madsen et al. (2024) proposed instruction-based self-consistency checks to measure faithfulness in generated explanations.
Another line of work by Wiegreffe et al. (2022) seeks to improve free text self-explanations with the help of human-written explanations that are included in the instruction. Similarly to Kunz and Kuhlmann (2024), self-explanations are evaluated on a variety of properties by the means of human annotation. They are found to be generally true, grammatical and factual Wiegreffe et al. (2022) and further selective, to contain illustrative examples and rarely subjective according to Kunz and Kuhlmann. For those two papers GPT-3 on CommonsenseQA/NLI and GPT-4 on the Alpaca dataset were analysed, respectively.
In our study, we consider human rationales as the ground truth for explaining a decision, against which we compare model self-explanations and post-hoc attributions.
@@
Our work focuses on a direct comparison of plausibility and faithfulness, using binary rationales and comparing them to post-hoc LRP attributions, which have been shown to faithfully reflect LLM predictions Ali et al. (2022); Achtibat et al. (2024). In particular, we extend our analysis beyond commonly used sentiment classification by considering more complex tasks such as forced labor detection from news articles and multilingual settings, further broadening the scope and applicability of our study.
@@ 3.1 Datasets
SST/mSST We use two different subsets from the Stanford Sentiment Treebank (SST2, Socher et al. 2013) for binary sentiment classification on movie reviews. The first subset (SST) contains 263 samples from the validation and test split from SST2 with an average sentence length of 18 tokens. Human rationale annotations have been published for that subset by Thorn Jakobsen et al. (2023) where each sample has been annotated by multiple annotators, 8 on average, who were recruited via Prolific. Annotators were first asked to classify the sample into one of three classes: positive, neutral or negative where none of the sentences was assigned neutral as a gold label. In a second step, annotators should choose the parts of the input that support their label choice. We select the rationale annotations with the correct labels from the first step for further analysis. We averaged the binary rationales across all annotators (with correct label classification) and set a threshold of $0.5$ (after averaging) for the token selection. We additionally analyse the rationale annotations collected by Jørgensen et al. (2022) on a subset of 250 samples from the validation set of SST2 (mSST). All samples were translated into Danish and Italian with an average sentence length of 15-17. Rationale annotation was carried out by 2 annotators per language (including English), who were native speakers with linguistic training. In contrast to the annotations collected by Thorn Jakobsen et al., the correct sentiment (positive or negative) was provided and the annotators were asked to select parts of the input that supported the gold label.
RaFoLa The authors of Mendez Guzman et al. (2022) published a Rationale-annotated corpus for Forced-Labour detection. This multi-class and multi-label dataset contains 989 English news articles that have been labelled and annotated according to 11 risk indicators defined by the International Labour Organization. Rationale annotations were carried out by two annotators who selected parts of the input to justify their label decision if they found evidence for any of the 11 labels. A subset of 100 articles was annotated by both annotators with a label agreement of $0.81$ (micro F1) and a rationale agreement of $0.73$ (intersection-over-union). The remaining articles were only annotated by one of the annotators. Each news article was assigned $1.2$ labels on average while $43\%$ were assigned with at least one label. For our analysis, we selected the 4 most frequent classes with occurrences between 117-256 out of the 989 articles. As we carry out zero-shot experiments on models that have not been fine-tuned on this task, we further convert this task into a binary classification task where we ask for a specific label once at a time. We provide the definition of the respective forced labour indicator as part of the instruction, see Figure 8 and Figure 9.
@@ 3.2 Rationale Extraction
For our experiments, we evaluate the following 4 instruction fine-tuned LLMs: Llama2-13B, Llama3.1-8B, Mistral-7B and quantized Mixtral-8x7B. 222meta-llama/Llama-2-13b-chat-hf333meta-llama/Meta-Llama-3.1-8B-Instruct444mistralai/Mistral-7B-Instruct-v0.3555mistralai/Mixtral-8x7B-Instruct-v0.1
In a first step, we ask the model to classify the given text into positive/negative for SST and into yes/no depending on evidence for a specific risk indicator for the RaFoLa dataset. If the model manages to generate the correct answer, we ask it to generate rationales based on the relevant provided context of the input. In case of RaFoLa, we follow the original data collection and only request rationales if the respective risk indicator is present. For the subsets in Italian and Danish, we have manually translated the prompts to the respective language with the help of native speakers.
Experimental Details The experiments are based on the transformers library. We set the repetition penalty to 1.0 and adjust the maximum length of generated text with respect to the task and expected output. We ensure reproducibility of our results by consistently using the same set of three seeds across our experiments and will release our code upon publication, including all parameters and the exact libraries used. All instructions with class definitions are presented in Appendix A.
3.3 Post-hoc Attribution
To extract input attribution scores, we use layer-wise relevance propagation (LRP); a widely used and state-of-the-art XAI method to compute feature attributions in LLMs Ali et al. (2022); Achtibat et al. (2024); Rezaei Jafari et al. (2024). Following the proposed propagation rules for Transformer models, we compute relevance scores for Llama and Mistral by backpropagating the logit for the correctly generated class token. While self-explanations and human annotations provide binary rationales, LRP assigns continuous scores to each token. To allow direct comparison, tokens are ranked based on their relevance scores, and the top-$k$ tokens–where $k$ is the number of rationales from the human annotation–are selected for further analysis.
@@ 3.4 Constraining Self-Explanations
Initial experiments show that without precise instructions, the model returns $80\%$ of the input tokens as rationales for SST where humans had annotated approximately $30\%$. This made comparisons difficult, so we chose to request a maximum number of tokens based on the number of annotated tokens by humans for each sample. Language models did not always follow this request but we could reduce the ratio of tokens to a comparable level with human annotations. For RaFoLa, this issue was less pronounced, as the input texts were much longer and humans annotated entire phrases. We thus decided not to include an upper bound for the RaFoLa rationales. We will discuss ratios and instruction following in the Appendix in Section C.
Table 1: Model accuracies for SST, multilingual SST, and RaFoLa, with highest scores shown in bold.
Data Acc. llama llama3 mistral mixtral
SST 0.88 0.98 0.98 0.98
mSST (EN) 0.85 0.98 0.98 1.00
mSST (DK) 0.84 0.87 0.97 0.97
mSST (IT) 0.88 0.95 0.97 0.99
RaFoLA #1 0.50 0.47 0.65 0.39
RaFoLA #2 0.58 0.58 0.58 0.48
RaFoLA #5 0.88 0.92 0.89 0.81
RaFoLA #8 0.90 0.90 0.90 0.90
4 Main Results
We first show and discuss the main results of task accuracy and pair-wise agreement between the different types of rationales, i.e., plausibility scores, before further analysing their faithfulness.
4.1 Task accuracy
Table 1 presents task accuracies for SST, multilingual SST (mSST), and RaFoLa. Accuracies for SST and mSST are generally high across models, while RaFoLa shows more variation and overall lower accuracies. We can assume that most models nowadays have seen the original English version of SST during training and are thus more familiar with this type of data. From the set of models we consider for this study, all have been pre-trained on English data while Llama3 and Mixtral also have been pre-trained on Italian, but none of the model officially supports Danish. Considering this difference in language exposure, our results show that all models are able to solve the sentiment classification task in Danish and Italian with accuracies comparable to English.
@@ 4.1 Task accuracy
#8 Physical and sexual violence asdf
Figure 2: Pair-wise comparison scores (Cohen’s Kappa) between rationales on SST and multilingual SST (upper) and RaFoLa (lower). We compare human-annotated, model-generated (self-explanations), and post-hoc rationales.
@@ 4.2 Plausibility
Following XAI literature, we assess plausibility of rationales to humans by considering human annotations as the ground truth and compute agreement to model rationales DeYoung et al. (2020). Here, we further include agreement scores between model-generated and post-hoc LRP rationales. We show pair-wise comparisons by calculating sample-wise Cohen’s Kappa scores between the binary scores and averaging across samples for different models.
Metric Cohen’s Kappa Cohen (1960) is a well-established method to measure inter-annotator agreement (IAA) between two annotators, in our case those are either the averaged human annotations or the two different types of model rationales (generated and post-hoc). We chose Kappa over F1 scores, which is also often used to evaluate IAA but comes with two obstacles. It is (i) driven by the imbalance of classes (here selected and not selected tokens) leading to a higher offset for annotations with a ratio of selected tokens closer to 0.5 and it (ii) does not consider randomness as a confounding factor. Cohen’s Kappa scores account for both issues, leading to overall lower but more robust scores than F1, which are also reported in the Appendix in Section B.
@@
SST
mSST      Danish
RaFoLa        article #1
article #8
Figure 3: Faithfulness evaluation for SST, mSST (Danish) and RaFoLa (articles #1 and #8). Model probability after masking the tokens extracted from human rationales, model self-explanation rationales and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
@@ 4.3 Faithfulness
Besides plausibility, faithfulness is the most commonly used evaluation approach to judge the quality of model explanations. Especially for feature attribution approaches, removal of most relevant features has been used to assess how faithful a feature subset is with respect to the model prediction, i.e., if removing a highly relevant subset will lead to a strong decrease of the prediction. We evaluate faithfulness by measuring the change in probability after masking the tokens as identified by the different rationales (human, model self-explanations and model post-hoc).
@@
Compared to human and self-explanation rationales, post-hoc attributions provide a relevance score for each token in the input prompt, requiring to binarize post-hoc attributions to allow for direct comparison as described in Section 3.3. Additionally, we included a baseline that randomly removes as many token as identified by humans. Our key results for SST and RaFoLa are summarized in Figure 3, with additional results across all languages and articles presented in Figure 15 in the Appendix. Average initial probability for the correct answer token is given as a dashed line for each model.
@@
We find that levels of faithfulness for humans, generated, and post-hoc explanations are comparable, with human explanations being overall as faithful as those generated by models or post-hoc methods. Self-generated model explanations can be more faithful than post-hoc ones, and the reverse case can also be true, depending on the task and model. Overall, they provide similarly faithful model rationales. We further investigate the limited impact of removing features on the class token probability in the post-hoc setting and found that the most relevant tokens are typically part of the provided task instruction, such as the class definition or question. Furthermore, contrastive explanations have been proposed to provide more task-specific attributions Yin and Neubig (2022), which we analyzed in Appendix E.2. We found no critical effect of contrastive explanations on the plausibility and faithfulness of model-based rationales.
[…]
Figure 4: Distribution of POS-tags in comparison to top-6 POS in human annotations for SST. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for model-generated, i.e., self-explanations (left) and post-hoc LRP (right) across all data splits and languages for Mistral.
#1 Abuse of vulnerability #8 Physical & sexual violence